I need help with this rational expression please:
I am supposed to build a rational expression into an equivalent
rational expression with the indicated denominator.
(-7yt)/3x=?/18xyt
Here is my try: (-7yt)/3x= -42yt/18xyt
Thanks

Answer by RAY100(1637)

Very good,,,just slightly off
.
Let the ? be the expression you would place in the numerator to fit.
.
(-7yt)/(3x) = ? / (18xyt)
.
lets cross multiply
.
(3x) * ? = (-7yt)(18xyt)
.
(3x)*? = -7*18 * x*y^2*t^2
.
divide thru by 3x
.
? = -42 y^2 t^2
.
to check
(-42y^2t^2) / (18xyt) = (-7 y t) / 3 x,,,,ok
.
Have a good day